OpenStudy (anonymous):

solve for y: y^2-81=0 this is solving factorable quadratic equations medal awarded

OpenStudy (cwrw238):

this is the difference of 2 squares a^2 - b^2 = (a + b)(a - b) = 0 now substitute y^2 for a^2 and 81 for b^2 what do you get?

OpenStudy (anonymous):

y^2-81=(y+81)(y-81)

OpenStudy (cwrw238):

not quite b^2 = 81 so b = 9 follow?

OpenStudy (cwrw238):

y^2 = 81 so y = 9

OpenStudy (anonymous):

yes okay so y^2-81=(y+9)(y-9)

OpenStudy (anonymous):

how is y^2=81

OpenStudy (cwrw238):

compare your equation with the general form: a^2 + b^2 y^2 + 81

OpenStudy (cwrw238):

ohh '-' not '+'

OpenStudy (anonymous):

so shouldnt it be -81 and not 81

OpenStudy (anonymous):

??

OpenStudy (anonymous):

@terenzreignz please help im all turned around now gah

terenzreignz (terenzreignz):

I'm terrible at teaching how to factor... as much as it pains me to say it... you need a second (third?) opinion :)

terenzreignz (terenzreignz):

although, with this one, what I can say to you is that if you a difference of two squares.. \[\Large a^2 - b^2\] then it factors into the sum and difference of their square roots.. \[\Large a^2 - b^2 = (a+b)(a-b)\]

terenzreignz (terenzreignz):

With that in mind.... \[\LARGE y ^2 - 81 = \color{blue}{y^2 - 9^2}\] Get to it, champ :)

terenzreignz (terenzreignz):

Wait a minute, you already finished this :/

OpenStudy (anonymous):

so it is -81 and -9

OpenStudy (anonymous):

?

terenzreignz (terenzreignz):

No... lol (y + 9)(y - 9) = 0 So either y + 9 = 0 or y - 9 = 0 You have to consider them both :)

OpenStudy (anonymous):

9 and -9

terenzreignz (terenzreignz):

That's right :D

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

lets do another ill post it

terenzreignz (terenzreignz):

Okay... look at this way... if the product of two numbers is 0... ab = 0 then at least one of them must have been zero, right? So, either a = 0 or b = 0 It's not that different with polynomials... (x+9)(x-9) = 0 Then either x+9 = 0 or x-9 = 0 It's that simple ;)